OpenStudy (anonymous):

the flag pole bend 6m from the ground. the depression of its angle is 30'5'. how tall is the pole?

OpenStudy (anonymous):

wat do u mean by bend

OpenStudy (anonymous):

cos30.5=a/6 a=cos30.5*6 a=5.17m